# MA-LMM: Memory-Augmented Large Multimodal Model for Long-Term Video Understanding

### v1

@@ 1 Introduction
Figure 1: (a) We propose the long-term memory bank to auto-regressively store and accumulate past video information, different from previous methods directly feeding the visual encoder’s outputs into the querying transformer. (b) GPU memory consumption v.s. video frame length of existing multimodal methods and MA-LMM during inference. Circle sizes represent the number of text tokens.
@@ 2 Related Work
Video-language models. Current image-language models such as Flamingo [22] and BLIP-2 [7, 9] can also support video inputs. They simply flattened the spatio-temporal features into 1D sequences and then fed them into the language models for video inputs. However, these approaches can not effectively capture the temporal dynamics of videos. Recent video-language models [11, 12, 13] attempt to address this by incorporating mechanisms to better understand these dynamics. For instance, Video-LLaMA [12] enhances BLIP-2 structure by adding an additional video querying transformer to explicitly model the temporal relationship. Similarly, building on LLaVA [8], Video-ChatGPT [21] simply average pools the frame-level features across spatial and temporal dimensions to generate video-level representation. ChatVideo [24] treats tracklets as the basic video elements and relies on several video foundation models to annotate these tracklets with textual descriptions. Then, it uses ChatGPT [4] to achieve various video understanding tasks. VideoChat [13] utilizes perception models to generate action and object annotations, which are then forwarded to LLMs for further reasoning. Despite these advancements, these models are primarily designed for short videos, as longer videos pose significant challenges due to language model context length and GPU memory restrictions.
@@ 3 Method
We introduce MA-LMM, a memory-augmented large multimodal model for long-term video understanding. Figure 2(a) illustrates the overview of our MA-LMM framework. Following similar practices of large multimodal models [7, 9, 8, 12], the overall model architecture can be divided into three parts: (1) visual feature extraction with a frozen visual encoder (Sec. 3.1), (2) long-term temporal modeling with a trainable querying transformer (Q-Former) to align the visual and text embedding spaces (Sec. 3.2), and (3) text decoding with a frozen large language model (Sec. 3.3).
@@ 3.1 Visual Feature Extraction
We propose to auto-regressively process video frames in an online manner. This design draws inspiration from the cognitive processes humans use to handle long-term visual information. Instead of concurrently processing extensive duration of signals, humans process them in a sequential manner, correlate current visual inputs with past memories for comprehension, and selectively retain salient information for subsequent reference [38]. Similarly, our MA-LMM processes video frames sequentially, dynamically associating new frame input with historical data stored in the long-term memory bank, ensuring that only discriminative information is conserved for later use. This selective retention facilitates a more sustainable and efficient approach to video understanding, which further allows the model to automatically support online video reasoning tasks.
@@ 3.2 Long-term Temporal Modeling
Memory Bank Compression. Given that our model directly stores past video information in the memory banks, the GPU memory cost and computational complexity increase linearly as the number of past video frames. This becomes particularly challenging for long videos, and thus it is essential to further compress the memory bank to a smaller size. A straightforward solution is to use a first-in-first-out queue to manage the temporal sequence. In this scenario, the features of the earliest timestep are removed if the current length of the memory bank exceeds a pre-defined limit. However, this approach inevitably leads to the loss of earlier historical information as new frames are added and old features are popped when the memory bank is full. To address this issue, we take advantage of the temporal redundancy in long videos and introduce a novel memory bank compression (MBC) technique. This method aggregates and compresses video information over time based on the similarity of adjacent features, while preserving early historical information. In this manner, repetitive information is consolidated in the memory bank, while discriminative features remain intact.
We apply the compression algorithm at each auto-regressive iteration if the current memory bank length is larger than the pre-defined threshold $M$. Formally, given the visual memory bank containing a list of $f_{1},f_{2},..,f_{M}],f_{t}\in\mathbb{R}^{P\times C}$, when a new frame feature $f_{M+1}$ comes in, we need to compress the memory bank by reducing the length by 1. At each spatial location $i$, we first calculate the cosine similarity between all the temporally adjacent tokens as
@@ 4.5 Visualization
In Figure 4, we provide a comprehensive visual comparison between MA-LMM and Video-LLaMA [12]. In the video question answering task, our MA-LMM exhibits superior memorization and recognition capabilities. Specifically, it can accurately memorize historical information and recognize fine-grained information, such as the color of the man with No.7, and precisely count the number of goalkeepers who appeared in the video. With the auto-regressive design, our model supports online reasoning directly. This capability is further exemplified in our experiments on off-the-shelf evaluations using custom questions. We can see that, our model can correctly anticipate the next step of the video ("egg will be cooked") and predict the correct recipe of the video ("scrambled egg"). More visualization examples are shown in the supplementary material.
@@ 5 Conclusion
We report additional ablation experiments in Sec. A. We also present more qualitative results on the video captioning task in Sec. B. And in Sec. C, we show more dataset-specific implementation details and hyper-parameters. Finally, we discuss some limitations and future works in Sec. D.
@@ Appendix A Additional Experiments
Inference time of different input frames In Figure 6, the inference time of MA-LMM increases linearly with respect to the frame lengths, due to its auto-regressive design of processing video frames sequentially. In contrast, directly concatenating frame-level features takes much longer time and higher GPU memory consumption, since it needs to process all video frames simultaneously.
@@ Appendix B More Qualitative Results
Appendix C Experiment Details
We show the details of hyper-parameters in the following table for different tasks and datasets. For all the experiments, we use a cosine learning rate decay. Table 11 shows the hyper-parameters for the long-term video understanding task. For the LVU dataset, we follow the same practice in [32, 33], we sample 100 frames of 1 fps for each video clip. For the Breakfast [42] and COIN [43], we uniformly sample 100 frames from the whole video. Table 12 shows the hyper-parameters on the MSRVTT-QA [48], MSVD-QA [48], and ActivityNet-QA [49] datasets for the video question answering task while Table 13 presents the hyperparameters on the MSRVTT [52], MSVD [53], YouCook2 [54] datasets for the video captioning tasks.
@@ Appendix C Experiment Details
Appendix D Limitation and Future Work

### v2

@@ 1 Introduction
Figure 1: (a) We propose the long-term memory bank to auto-regressively store and accumulate past video information, different from previous methods directly feeding the visual encoder’s outputs into the querying transformer. (b) GPU memory and token number v.s. video frame length of multimodal methods and MA-LMM during inference. Circle sizes represent the number of text tokens.
@@ 2 Related Work
Video-language models. Previous image-language models such as Flamingo [22] and BLIP-2 [7, 9] can also support video inputs. They simply flattened the spatio-temporal features into 1D sequences and then fed them into the language models for video inputs. However, these approaches can not effectively capture the temporal dynamics of videos. Based on this motivation, Video-LLaMA [12] enhances BLIP-2 structure by adding an additional video querying transformer to explicitly model the temporal relationship. Similarly, building on LLaVA [8], Video-ChatGPT [21] simply average pools the frame-level features across spatial and temporal dimensions to generate video-level representation. VideoChat [13] utilizes perception models to generate action and object annotations, which are then forwarded to LLMs for further reasoning. Despite the advancements, these models are primarily designed for short videos. Inspired by the Token Merging [24] which averages highly similar tokens to reduce the computation cost, we propose an extension of this idea to video data, specifically along the temporal axis. This extension aims to mitigate the challenges posed by extensive token numbers and computational cost associated with processing long video inputs. Several concurrent works [25, 26, 27] have also explored similar strategies of merging akin tokens for video inputs. Please refer to the supplementary material for more detailed discussions.
@@ 3 Method
We introduce MA-LMM, a memory-augmented large multimodal model for long-term video understanding. Instead of processing more frames simultaneously as most video understanding methods [42, 31, 43, 44, 45, 46, 47, 48, 49], we propose to auto-regressively process video frames in an online manner, which draws inspiration from the online processing fashion with long-term memory design presented in MeMViT [41]. Figure 2(a) illustrates the overview of our MA-LMM framework. Following similar practices of large multimodal models [7, 9, 8, 12], the overall model architecture can be divided into three parts: (1) visual feature extraction with a frozen visual encoder (Sec. 3.1), (2) long-term temporal modeling with a trainable querying transformer (Q-Former) to align the visual and text embedding spaces (Sec. 3.2), and (3) text decoding with a frozen large language model (Sec. 3.3).
@@ 3.1 Visual Feature Extraction
This design draws inspiration from the cognitive processes humans use to handle long-term visual information. Instead of concurrently processing extensive duration of signals, humans process them in a sequential manner, correlate current visual inputs with past memories for comprehension, and selectively retain salient information for subsequent reference [41]. Similarly, our MA-LMM processes video frames sequentially, dynamically associating new frame input with historical data stored in the long-term memory bank, ensuring that only discriminative information is conserved for later use. This selective retention facilitates a more sustainable and efficient approach to video understanding, which further allows the model to automatically support online video reasoning tasks.
@@ 3.2 Long-term Temporal Modeling
Memory Bank Compression. Given that our model directly stores past video information in the memory banks, the GPU memory and computational cost increase linearly as the number of past video frames. This becomes particularly challenging for long videos, and thus it is essential to further compress the memory bank to a smaller size. One conventional approach to managing temporal sequences involves employing a first-in-first-out queue. Here, features from the earliest time step are removed when the memory bank reaches a pre-defined limit, a strategy utilized in MeMViT [41]. However, it results in the loss of earlier historical information as new frames are added and old features are popped to maintain memory bank capacity. Alternatively, MeMViT employs learnable pooling operators to compress the spatio-temporal size of stored feature in the memory bank, albeit at the cost of introducing additional trainable parameters.
Drawing inspiration from the effectiveness of token merging and pruning techniques showcased in works such as [24, 50, 51, 52], we introduce a novel Memory Bank Compression (MBC) technique to exploit temporal redundancies inherent in videos. Our proposed method aggregates and compresses video information over time by leveraging the similarity between adjacent features, thereby retaining early historical information. This approach effectively compresses repetitive information within the memory bank while preserving discriminative features. Notably, several concurrent works [25, 26, 27] have similarly embraced the token merging strategies to reduce video redundancies.
Same as MeMViT [41], which applies feature compression at each iteration, our method applies the compression algorithm at each auto-regressive step if the current length of the memory bank exceeds the predefined threshold $M$. Formally, given the visual memory bank containing a list of $f_{1},f_{2},..,f_{M}],f_{t}\in\mathbb{R}^{P\times C}$, when a new frame feature $f_{M+1}$ comes in, we need to compress the memory bank by reducing the length by 1. At each spatial location $i$, we first calculate the cosine similarity between all the temporally adjacent tokens as
@@ 4.5 Visualization
In Figure 4, we provide a comprehensive visual comparison between MA-LMM and Video-LLaMA [12]. In the video question answering task, MA-LMM exhibits superior memorization and recognition capabilities. Specifically, it can accurately memorize historical information and recognize fine-grained information, such as the color of the man with No.7, and precisely count the number of goalkeepers who appeared in the video. With the auto-regressive design, our model supports online reasoning directly. This capability is further exemplified in our experiments on off-the-shelf evaluations using custom questions. MA-LMM can correctly anticipate the next step of the video ("egg will be cooked") and predict the correct recipe ("scrambled egg"). More visualization examples are shown in the supplementary material.
@@ 5 Conclusion
We present additional ablation experiments in Section A and further qualitative results for the video captioning task in Section B. Next, in Section C, we discuss the relations to concurrent works [25, 26, 27] in details. And in Sec. D, we show more dataset-specific implementation details and hyper-parameter settings. Finally, we address some limitations and outline directions for future research in Section E.
@@ Appendix A Additional Experiments
Inference time v.s. video frame lengths. In Figure 6, the inference time of MA-LMM increases linearly with respect to the frame lengths, due to its auto-regressive design of processing video frames sequentially. In contrast, directly concatenating frame-level features takes much longer time and higher GPU memory consumption, since it needs to process all video frames simultaneously.
@@ Appendix B More Qualitative Results
Appendix C Relations to Concurrent Works
In this section, we compare and discuss the relations between our MA-LMM with the concurrent works including TESTA [25], MovieChat [26] and Chat-UniVi [27]. All of these methods focus on utilizing the idea of token merging [24] to reduce video redundancies.
Temporal Modeling. Temporal modeling across these methodologies falls into three categories. Chat-UniVi [27] directly feed visual tokens into large language models (LLMs) without explicit temporal modeling, utilizing LLMs’ inherent sequence processing for video understanding. In contrast, TESTA [25] and MovieChat [26] employ global self-attention; TESTA captures interactions along spatial and temporal dimensions, whereas MovieChat processes long videos in segments, compresses these into short-term memories, then concatenates and models global temporal interactions using a video Q-Former. Differently, our MA-LMM adopts causal self-attention, restricting each frame’s feature access to prior video information only. Such a design naturally endows our MA-LMM with the capability to support online video applications in robotics, AR/VR, and video streaming.
Token Merging Application. Building on the token merging [24] strategy, four methodologies have adopted and modified this approach to reduce video data redundancy. Each uses the core concept of merging similar tokens but differs in implementation. TESTA [25] utilizes a cascaded module for spatial and temporal aggregation, progressively shortening video length and decreasing tokens per frame. In contrast, Chat-UniVi’s [27] modules operate in parallel, merging tokens across both dimensions before LLM reasoning. MovieChat [26] employs a selective strategy to merge similar adjacent frames, reducing the number of video frames. Similarly, our MA-LMM conducts token merging along the temporal dimension to condense video length but at a more fine-grained spatial level. It independently compresses visual and query tokens across different spatial areas, enhancing performance as evidenced in Table 10.
Based Model. Both TESTA [25] and Moviechat [26] are built upon the video-based multimodal model. TESTA integrates TimeSFormer [47] as its video encoder, facilitating long-range video modeling. Meanwhile, MovieChat adopts the Video-LLaMA [12] framework, combining an image Q-Former with a video Q-Former to effectively manage long-term temporal relationships. On the contrary, another group involves adapting image-based multimodal models for video understanding. Chat-UniVi [27] leverages the LLaVA [8] architecture, feeding concatenated visual tokens along the temporal axis into LLMs. Our MA-LMM builds on InstructBLIP [9] as a plug-and-play module that significantly boosts long-term temporal modeling. Demonstrated in Table 9, our memory bank module greatly excels over InstructBLIP under the off-the-shelf setting without video-specific pre-training or introducing additional parameters.
Memory Bank Design. The integration of memory banks to enhance long-term video understanding has been thoroughly explored [39, 87, 88, 38, 41]. Building on these studies, MovieChat [26] and our MA-LMM both employ memory bank designs. MovieChat primarily uses memory banks to consolidate raw and static visual features. In contrast, our MA-LMM innovates with an additional query memory bank that captures dynamic memory, reflecting the evolving understanding of past video frames. The effectiveness of our query memory bank is evidenced in Table 9.
Appendix D Experiment Details
We build our MA-LMM on top of InstructBlip [9], following the codebase [89]. We show the details of hyper-parameters in the following table for different tasks and datasets. For all the experiments, we use a cosine learning rate decay. Table 11 shows the hyper-parameters for the long-term video understanding task. For the LVU dataset, we follow the same practice in [35, 36], we sample 100 frames of 1 fps for each video clip. For the Breakfast [56] and COIN [57], we uniformly sample 100 frames from the whole video. Table 12 shows the hyper-parameters on the MSRVTT-QA [62], MSVD-QA [62], and ActivityNet-QA [63] datasets for the video question answering task while Table 13 presents the hyperparameters on the MSRVTT [66], MSVD [67], YouCook2 [68] datasets for video captioning.
Appendix E Limitation and Future Work